# Large Language Models for Multilingual Vulnerability Detection: How Far Are We?

### v1

@@ 4.1. RQ1: How effective are PLMs and LLMs in detecting multilingual vulnerabilities at the function-level?
Observation 2: LLMs exhibit limitations in detecting function-level multilingual vulnerability through zero-shot or few-shot prompting strategies alone. In evaluating the effectiveness of LLMs for function-level multilingual vulnerability detection under zero-shot and few-shot prompting settings, Accuracy emerges as a critical metric due to the balanced nature of the dataset. Under zero-shot prompting setting, the average Accuracies of LLMs across seven programming languages were 0.5005 for DeepSeek-Coder, 0.4877 for Code Llama, 0.5005 for Llama 3, 0.4758 for GPT-3.5-Turbo, and 0.4985 for GPT-4. Specifically, DeepSeek-Coder achieves the highest Accuracy (0.5005), closely followed by Llama 3 (0.5005) and Code Llama (0.4877). Although DeepSeek-Coder exhibits the highest Accuracy, it achieves the lowest Recall (0.1796), leading to a poor F1-score (0.2656). This suggests a tendency to overpredict non-vulnerabilities. In contrast, Code Llama stands out with the highest Recall (0.9156). However, its relatively lower Precision (0.4950) results in a moderate F1-score (0.6426), indicating a tendency to overpredict vulnerabilities, leading to a higher FPR (0.9454) GPT-3.5-Turbo and GPT-4o demonstrate lower Recall but more balanced Precision-Recall trade-offs, though with modest Accuracy scores (0.4827 and 0.4985, respectively).
@@
Under the few-shot prompting setting, this strategy does not consistently improve accuracy. In fact, most LLMs show decreased performance compared to zero-shot prompting, with GPT-4o being the only exception. Specifically, DeepSeek-Coder drops sharply to 0.3870 Accuracy, Code Llama to 0.4763, Llama 3 to 0.4299, and GPT-3.5-Turbo to 0.3618, indicating that providing few examples may confuse rather than help the models. In this case, GPT-4o is the only exception, slightly improving to 0.5035 Accuracy, though its Recall (0.4347) and F1-score (0.4683) remain relatively low. These results highlight that zero-shot prompting yields more stable Accuracy for LLMs, and few-shot prompting does not guarantee improvement. Overall, LLMs using only zero-shot or few-shot prompting show limited effectiveness in function-level multilingual vulnerability detection, emphasizing the need for further adaptation to the task domain.
((a))
((b))
@@
Figure 4. Performance of PLMs and LLMs across seven programming languages in function-level multilingual vulnerability detection (y-axis: Accuracy)
@@
Observation 3: Instruction tuning is critical to improve the LLM’s effectiveness in detecting function-level multilingual vulnerability. Instruction tuning is able to enhance the function-level multilingual vulnerability detection capabilities of LLMs, especially when paired with prompting strategies. In the instruction tuning with zero-shot prompting setting, GPT-4o achieves the highest Accuracy (0.5874), outperforming GPT-3.5-Turbo (0.5454), Code Llama (0.4783), Llama 3 (0.5192), and DeepSeek-Coder (0.5039). Notably, most of the instruction-tuned LLMs show improvement in Accuracy, Precision, and F1-scores, compared to their non-instruction-tuned counterparts. For instance, the most significant accuracy improvements were seen in GPT-4o (17.83%) and GPT-3.5-Turbo (14.63%). These results suggest that instruction tuning enables better generalization and contextual understanding of multilingual vulnerability patterns, even without in-context examples.
@@
When few-shot prompting was combined with instruction tuning, most LLMs exhibited the same tendency as with few-shot prompting alone, i.e., LLMs’ performance deteriorated, except for GPT-4o. In this case, GPT-4o achieves a remarkable Accuracy of 0.7196, the highest across all LLMs with different strategies, along with the highest Precision (0.7454) and F1-score (0.7069). Additionally, GPT-4o achieves a relatively balanced ratio between its FNR of 0.3278 and FPR of 0.2324, indicating its ability to effectively detect vulnerable functions while maintaining accurate predictions for clean functions. This suggests that GPT-4 has strong discriminative capabilities and robustness in detecting function-level multilingual vulnerability. Compared to the best-performing PLM, CodeT5P, GPT-4o exhibits a significant advantage in Accuracy and Precision, while maintaining a competitive F1-score. Although CodeT5P shows higher Recall (0.9529 vs. 0.6722), CodeT5P’s FPR (0.7496) is also significantly higher than GPT-4o’s FPR (0.2324). In contrast, GPT-4o’s balanced trade-off across all metrics, particularly its much lower harmonic mean between FPR and FNR ($\frac{0.2324+0.3278}{2}$ < $\frac{0.7496+0.0471}{2}$), making it more practical in real-world applications. This comparison underscores the growing potential of instruction-tuned LLMs, especially GPT-4o, as powerful and adaptable techniques for function-level multilingual vulnerability detection when paired with minimal task-specific examples.
@@
4.2. RQ2: How effective are PLMs and LLMs in detecting multilingual vulnerabilities at the line level?
@@
To evaluate the effectiveness of the studied PLMs and LLMs in detecting line-level multilingual vulnerability, we used common classification metrics: Accuracy, Precision, Recall, F1-score, FPR, and FNR, as detailed in 3.4. Since our task involves classifying each line of code within a function, we faced a challenge with imbalanced data, i.e., most code lines are non-vulnerable, creating an uneven distribution in our line-level dataset. For this reason, we chose F1-score as our primary metric.
@@
Table 4. Performance of PLMs and LLMs for line-level multilingual vulnerability detection
@@
Techniques Accuracy Recall Precision F1-score FPR FNR Dummy Classifier ${DummyClf}_{vul}$ 0.0267 1.0000 0.0267 0.0520 1.0000 0.0000 ${DummyClf}_{clean}$ 0.9733 0.0000 0.0000 0.0000 0.0000 1.0000 PLMs CodeBERT 0.9203 0.8822 0.2353 0.3715 0.0787 0.1177 UniXcoder 0.8880 0.8879 0.1787 0.2975 0.1120 0.1121 CodeT5 0.9472 0.8119 0.3123 0.4511 0.0491 0.1881 CodeT5P 0.9563 0.7673 0.3536 0.4841 0.0385 0.2327 LLMs (Zero-Shot Prompting) DeepSeek-Coder 0.9536 0.3002 0.2246 0.2570 0.0284 0.6997 Code Llama 0.9589 0.2237 0.2272 0.2254 0.0208 0.7763 Llama 3 0.9307 0.4847 0.1891 0.2720 0.0570 0.5153 GPT-3.5-Turbo 0.9557 0.3215 0.2475 0.2797 0.0229 0.8393 GPT-4o 0.9537 0.4286 0.2700 0.3313 0.0318 0.5714 LLMs (Few-Shot Prompting) DeepSeek-Coder 0.9521 0.3493 0.2344 0.2805 0.0313 0.6507 Code Llama 0.9666 0.3018 0.3536 0.3257 0.0151 0.6982 Llama 3 0.8980 0.7509 0.1738 0.2822 0.0980 0.2491 GPT-3.5-Turbo 0.9774 0.4372 0.6057 0.5078 0.0078 0.5628 GPT-4o 0.9802 0.5689 0.6461 0.6050 0.0086 0.4311 LLMs (Instruction Tuning + Zero-Shot Prompting) DeepSeek-Coder 0.9404 0.2155 0.1297 0.1619 0.0397 0.7845 Code Llama 0.8771 0.5902 0.1234 0.2042 0.1150 0.4098 Llama 3 0.9415 0.3403 0.1819 0.2371 0.0420 0.6597 GPT-3.5-Turbo 0.9578 0.5186 0.3211 0.3966 0.0301 0.4814 GPT-4o 0.9729 0.4241 0.4910 0.4551 0.0121 0.5759 LLMs (Instruction Tuning + Few-Shot Prompting) DeepSeek-Coder 0.7149 0.2851 0.0278 0.0507 0.2733 0.7149 Code Llama 0.8688 0.6221 0.1207 0.2022 0.1243 0.3779 Llama 3 0.9220 0.4720 0.1647 0.2442 0.0657 0.5280 GPT-3.5-Turbo 0.9716 0.5804 0.4735 0.5215 0.0177 0.4196 GPT-4o 0.9830 0.6307 0.7012 0.6641 0.0074 0.3693
@@
Results. Table 4 shows the comparison results among PLMs and LLMs in terms of average Accuracy, Recall, Precision, F1-score, FPR, and FNR in detecting line-level multilingual vulnerability. Within each metric, the values in bold indicate the model that exhibits the best performance among all PLMs and LLMs. Figure 5 presents the performance of PLMs, LLMs with zero-shot prompting, and LLMs with instruction tuning and few-shot prompting in seven programming languages in terms of F1-score.
@@
Observation 5: CodeT5P is the best-performing model in identifying line-level multilingual vulnerability among PLMs. Among the evaluated PLMs, CodeT5P demonstrates the strongest overall performance in detecting line-level vulnerabilities, achieving the highest F1-score of 0.4841. This comparatively high F1-score primarily stems from its balanced approach, reflected in a moderate Recall of 0.7673 and Precision of 0.3536. Additionally, CodeT5P balanced the trade-off between predicting false positives and false negatives, achieving the lowest FPR (0.0385) among studied PLMs. Although UniXCoder achieves higher Recall (0.8879), indicating sensitivity in identifying vulnerable lines, its significantly lower Precision (0.1787) and higher FPR (0.1120) severely impacts its overall performance, yielding a much lower F1-score of 0.2975. CodeBert exhibits the modest performance among PLMs, with an F1-score of 0.3715, despite having relatively high Recall (0.8822). Thus, among traditional PLMs, CodeT5P emerges as the most viable choice for line-level vulnerability detection, effectively balancing FPR and FNR to mitigate the challenges of the imbalanced dataset.
@@
Observation 6: While using prompting alone, few-shot prompting can improve the effectiveness of LLMs in detecting line-level multilingual vulnerability. Under zero-shot prompting, all evaluated LLMs show limited practical effectiveness despite achieving notably high Accuracy scores. Code Llama attains the highest Accuracy (0.9589); however, it exhibits an extremely low F1-score (0.2254), reflecting significant weaknesses in Precision (0.2272), Recall (0.2237), and FNR (0.7763). Similarly, models like GPT-4o and GPT-3.5-Turbo display relatively high Accuracies (0.9537 and 0.9557, respectively) but similarly low F1-scores (0.3313 and 0.2797) and high FNR (0.5714 and 0.8393). The consistently low Recall/Precision and high FNR across zero-shot prompting indicate a clear inability of LLMs to effectively detect vulnerable lines without further guidance or contextual clues, underscoring the insufficiency of using LLMs in isolation under zero-shot conditions for multilingual line-level vulnerability detection.
@@
Few-shot prompting improves LLMs’ performance through in-context examples. GPT-4o demonstrates the best performance, surpassing all other PLMs with a significantly higher F1-score of 0.6050, supported by strong Precision (0.6461) and moderate Recall (0.5689). Compared to the zero-shot prompting setting, GPT-4o’s FPR significantly decreases from 0.0318 to 0.0086, while its FNR also reduces from 0.5714 to 0.4311. GPT-3.5-Turbo also shows notable improvement, with its F1-score rising to 0.5078. This improvement is primarily driven by enhanced Precision (0.6057), although Recall remains limited (0.4372). These results demonstrate that few-shot prompting alone, without instruction tuning, can partially address the issue of imbalanced labels (where vulnerable lines are fewer than non-vulnerable lines) to improve LLMs’ performance in detecting multilingual vulnerabilities at the line-level.
((a))
((b))
@@
Figure 5. Performance of PLMs and LLMs across seven programming languages in line-level multilingual vulnerability detection (y-axis: F1-score)
@@
Observation 7: Instruction tuning is also beneficial for improving LLMs in detecting multilingual vulnerabilities at the line-level. Finally, the addition of instruction tuning combined with few-shot prompting significantly enhances LLM performance, clearly outperforming all other prompting strategies. GPT-4o, in particular, achieves the highest overall performance with an impressive F1-score of 0.6641. This high F1-score results from a robust Precision of 0.7012 and substantial Recall of 0.6307, effectively handling the challenge posed by dataset imbalance. Additionally, GPT-4o reaches the lowest FPR of 0.0074 and the lowest FNR of 0.3693 among the studied LLMs (except for Llama 3 with few-shot prompting). This indicates an excellent balance between correctly identifying vulnerable and clean lines. GPT-3.5-Turbo also shows notable improvements under this condition, achieving an F1-score of 0.5215, demonstrating the broader positive impact of instruction tuning combined with few-shot prompting. While using instruction tuning with few-shot prompting, GPT-4o and GPT-3.5-Turbo significantly outperform the dummy classifier in F1-score and Precision. This demonstrates that their predictions are substantially more accurate than random guesses for line-level detection.
@@
When comparing GPT-4o with the best-performing PLM, CodeT5P (F1-score: 0.4841), GPT-4o’s superiority becomes even more evident. GPT-4o notably surpasses CodeT5P not only in F1-score (0.6641 vs. 0.4841) but also significantly in Precision (0.7012 vs. 0.3536). Although CodeT5P achieves slightly higher Recall (0.7673 vs. GPT-4o’s 0.6307), its lower Precision substantially limits its practical utility, as it generates many more false-positive predictions. Supporting evidence comes from the FPR and FNR metrics. GPT-4o achieves a much lower FPR (0.0074 vs. CodeT5P’s 0.0385), though it has a higher FNR (0.3693 vs. CodeT5P’s 0.2327).
@@
4.3. RQ3: What are the strengths and weaknesses of the PLMs and LLMs in multilingual vulnerability detection?
@@
  • The unique correct and incorrect detection. Certain PLMs and LLMs may be prone to detecting more vulnerabilities. Therefore, we further explored the orthogonality between the studied PLMs and LLMs by using Venn diagrams to assess the unique correct detection and incorrect detection. To generate the Venn diagrams, we first categorized correct/incorrect detections by PLMs and LLMs into distinct sets. We then compared these sets to analyze their overlap and unique characteristics. This analysis revealed which correct/incorrect detections were made by PLMs alone, by LLMs alone, and by both models together.
@@
Figure 6. The unique correct/incorrect detection of function-level multilingual vulnerability in PLMs and LLMs
Figure 7. The unique correct/incorrect detection of line-level multilingual vulnerability in PLMs and LLMs
@@
Table 5. The percentage of correct detection for function-level multilingual vulnerability detection using two representative techniques across the Top 25 most dangerous CWE-IDs in 2023

### v2

@@ 3.1 RQ1: How effective are PLMs and LLMs in detecting multilingual vulnerabilities at the function-level?
Observation 2: LLMs exhibit limitations in detecting function-level multilingual vulnerability through zero-shot or few-shot prompting strategies alone. In evaluating the effectiveness of LLMs for function-level multilingual vulnerability detection under zero-shot and few-shot prompting settings, Accuracy emerges as a critical metric due to the balanced nature of the dataset. Under zero-shot prompting setting, the average Accuracies of LLMs across seven programming languages were 0.5005 for DeepSeek-Coder, 0.4877 for Code Llama, 0.5005 for Llama 3, 0.4758 for GPT-3.5-Turbo, and 0.4985 for GPT-4. Specifically, DeepSeek-Coder achieves the highest Accuracy (0.5005), closely followed by Llama 3 (0.5005) and Code Llama (0.4877). Although DeepSeek-Coder exhibits the highest Accuracy, it achieves the lowest Recall (0.1796), leading to a poor F1-score (0.2656). This suggests a tendency to overpredict non-vulnerabilities. In contrast, Code Llama stands out with the highest Recall (0.9156). However, its relatively lower Precision (0.4950) results in a moderate F1-score (0.6426), indicating a tendency to overpredict vulnerabilities, leading to a higher FPR (0.9454). The MCC values for these models are either near zero or negative (e.g., -0.0585 for Code Llama and -0.0375 for GPT-3.5-Turbo), with AUC scores mostly below 0.5, indicating that they struggle to perform better than a random baseline in this setting. GPT-3.5-Turbo and GPT-4o demonstrate lower Recall but more balanced Precision-Recall trade-offs, though with modest Accuracy scores (0.4827 and 0.4985, respectively). Furthermore, we examined GRACE, which integrates AST structural information through zero-shot prompting. As shown in Table 3, GRACE achieves the highest Accuracy (0.5061) and Precision (0.5323) among all zero-shot baselines, outperforming general-purpose LLMs. However, this structural constraint induces a highly conservative prediction strategy. GRACE records the lowest Recall (0.1673) and F1-score (0.2546), along with the lowest FPR (0.1495). While explicit syntax information helps filter out benign code, reducing false alarms, it significantly hinders the model’s ability to generalize and identify diverse vulnerability patterns. This results in a high FNR (0.8327) and minimal MCC (0.0244) and AUC (0.5089).
@@
Under the few-shot prompting setting, this strategy does not consistently improve performance. In fact, most LLMs show decreased performance compared to zero-shot prompting, with GPT-4o being the only exception. Specifically, DeepSeek-Coder drops sharply to 0.3870 Accuracy, Code Llama to 0.4763, Llama 3 to 0.4299, and GPT-3.5-Turbo to 0.3618, indicating that providing few examples may confuse rather than help the models. In this case, GPT-4o is the only exception, slightly improving to 0.5035 Accuracy, though its Recall (0.4347) and F1-score (0.4683) remain relatively low, and its MCC (0.0078) and AUC (0.5039) show negligible improvement over the zero-shot baseline. These results highlight that zero-shot prompting yields more stable Accuracy for LLMs, and few-shot prompting does not guarantee improvement. Overall, LLMs using only zero-shot or few-shot prompting show limited effectiveness in function-level multilingual vulnerability detection, emphasizing the need for further adaptation to the task domain.
@@
Figure 4: Performance of PLMs and LLMs across seven programming languages in function-level multilingual vulnerability detection (y-axis: Accuracy)
@@
Observation 3: Instruction tuning is critical to improve the LLM’s effectiveness in detecting function-level multilingual vulnerability. Instruction tuning is able to enhance the function-level multilingual vulnerability detection capabilities of LLMs, especially when paired with prompting strategies. In the instruction tuning with zero-shot prompting setting, GPT-4o achieves the highest Accuracy (0.5874), outperforming GPT-3.5-Turbo (0.5454), Code Llama (0.4783), Llama 3 (0.5192), and DeepSeek-Coder (0.5039). Notably, most of the instruction-tuned LLMs show improvement in Accuracy, Precision, and F1-scores, compared to their non-instruction-tuned counterparts. Specifically, GPT-4o improves its MCC to 0.1820 and AUC to 0.5881, demonstrating that instruction tuning effectively boosts the model’s discriminative capabilities even without examples. For instance, the most significant Accuracy improvements were seen in GPT-4o (17.83%) and GPT-3.5-Turbo (14.63%). These results suggest that instruction tuning enables better generalization and contextual understanding of multilingual vulnerability patterns, even without in-context examples.
@@
When few-shot prompting was combined with instruction tuning, most LLMs exhibited the same tendency as with few-shot prompting alone, i.e., LLMs’ performance deteriorated, except for GPT-4o. In this case, GPT-4o achieves a remarkable Accuracy of 0.7196, the highest across all LLMs with different strategies, along with the highest Precision (0.7454) and F1-score (0.7069). Additionally, GPT-4o achieves a relatively balanced ratio between its FNR of 0.3278 and FPR of 0.2324, indicating its ability to effectively detect vulnerable functions while maintaining accurate predictions for clean functions. This configuration also produces the highest overall MCC (0.4360) and AUC (0.7169), signifying a significant leap in robustness over other approaches. This suggests that GPT-4 has strong discriminative capabilities and robustness in detecting function-level multilingual vulnerability. Compared to the best-performing PLM, CodeT5P, GPT-4o exhibits a significant advantage in Accuracy and Precision, while maintaining a competitive F1-score. Although CodeT5P shows higher Recall (0.9529 vs. 0.6722), CodeT5P’s FPR (0.7496) is also significantly higher than GPT-4o’s FPR (0.2324). In contrast, GPT-4o’s balanced trade-off across all metrics, particularly its much lower harmonic mean between FPR and FNR ($\frac{0.2324+0.3278}{2}$ < $\frac{0.7496+0.0471}{2}$), making it more practical in real-world applications. This comparison underscores the growing potential of instruction-tuned LLMs, especially GPT-4o, as powerful and adaptable techniques for function-level multilingual vulnerability detection when paired with minimal task-specific examples.
@@
3.2 RQ2: How effective are PLMs and LLMs in detecting multilingual vulnerabilities at the line level?
@@
To evaluate the effectiveness of the studied PLMs and LLMs in detecting line-level multilingual vulnerability, we used common classification metrics: Accuracy, Precision, Recall, F1-score, FPR, FNR, MCC, and AUC, as detailed in 2.4. Since our task involves classifying each line of code within a function, we faced a challenge with imbalanced data, i.e., most code lines are non-vulnerable, creating an uneven distribution in our line-level dataset. For this reason, we chose F1-score as our primary metric.
@@
Table 4: Performance of PLMs and LLMs for line-level multilingual vulnerability detection
@@
Techniques Accuracy Recall Precision F1-score FPR FNR MCC AUC Dummy Classifiers ${DummyClf}_{vul}$ 0.0267 1.0000 0.0267 0.0520 1.0000 0.0000 0.0000 0.5000 ${DummyClf}_{clean}$ 0.9733 0.0000 0.0000 0.0000 0.0000 1.0000 0.0000 0.5000 PLMs CodeBERT 0.9203 0.8822 0.2353 0.3715 0.0787 0.1177 0.4316 0.9018 UniXcoder 0.8880 0.8879 0.1787 0.2975 0.1120 0.1121 0.3688 0.8880 CodeT5 0.9472 0.8119 0.3123 0.4511 0.0491 0.1881 0.4838 0.8814 CodeT5P 0.9563 0.7673 0.3536 0.4841 0.0385 0.2327 0.5029 0.8644 LLMs (Zero-Shot Prompting) DeepSeek-Coder 0.9536 0.3002 0.2246 0.2570 0.0284 0.6997 0.2362 0.6359 Code Llama 0.9589 0.2237 0.2272 0.2254 0.0208 0.7763 0.2043 0.6014 Llama 3 0.9307 0.4847 0.1891 0.2720 0.0570 0.5153 0.2730 0.7138 GPT-3.5-Turbo 0.9557 0.3215 0.2475 0.2797 0.0268 0.6785 0.2596 0.6473 GPT-4o 0.9537 0.4286 0.2700 0.3313 0.0318 0.5714 0.3175 0.6984 LLMs (Few-Shot Prompting) DeepSeek-Coder 0.9521 0.3493 0.2344 0.2805 0.0313 0.6507 0.2622 0.6590 Code Llama 0.9666 0.3018 0.3536 0.3257 0.0151 0.6982 0.3097 0.6433 Llama 3 0.8980 0.7509 0.1738 0.2822 0.0980 0.2491 0.3295 0.8265 GPT-3.5-Turbo 0.9774 0.4372 0.6057 0.5078 0.0078 0.5628 0.5035 0.7147 GPT-4o 0.9802 0.5689 0.6461 0.6050 0.0086 0.4311 0.5962 0.7802 LLMs (Instruction Tuning + Zero-Shot Prompting) DeepSeek-Coder 0.9404 0.2155 0.1297 0.1619 0.0397 0.7845 0.1376 0.5879 Code Llama 0.8771 0.5902 0.1234 0.2042 0.1150 0.4098 0.2296 0.7376 Llama 3 0.9415 0.3403 0.1819 0.2371 0.0420 0.6597 0.2207 0.6491 GPT-3.5-Turbo 0.9578 0.5186 0.3211 0.3966 0.0301 0.4814 0.3877 0.7443 GPT-4o 0.9729 0.4241 0.4910 0.4551 0.0121 0.5759 0.4425 0.7060 LLMs (Instruction Tuning + Few-Shot Prompting) DeepSeek-Coder 0.7149 0.2851 0.0278 0.0507 0.2733 0.7149 0.0042 0.5059 Code Llama 0.8688 0.6221 0.1207 0.2022 0.1243 0.3779 0.2329 0.7488 Llama 3 0.9220 0.4720 0.1647 0.2442 0.0657 0.5280 0.2464 0.7031 GPT-3.5-Turbo 0.9716 0.5804 0.4735 0.5215 0.0177 0.4196 0.5098 0.7813 GPT-4o 0.9830 0.6307 0.7012 0.6641 0.0074 0.3693 0.6563 0.8116
@@
Results. Table 4 shows the comparison results among PLMs and LLMs in terms of average Accuracy, Recall, Precision, F1-score, FPR, FNR, MCC, and AUC in detecting line-level multilingual vulnerability. Within each metric, the values in bold indicate the model that exhibits the best performance among all PLMs and LLMs. Figure 5 presents the performance of PLMs, LLMs with zero-shot prompting, and LLMs with instruction tuning and few-shot prompting in seven programming languages in terms of F1-score.
@@
Observation 5: CodeT5P is the best-performing model in identifying line-level multilingual vulnerability among PLMs. Among the evaluated PLMs, CodeT5P demonstrates the strongest overall performance in detecting line-level vulnerabilities, achieving the highest F1-score of 0.4841. This comparatively high F1-score primarily stems from its balanced approach, reflected in a moderate Recall of 0.7673 and Precision of 0.3536. Furthermore, CodeT5P achieves the highest MCC (0.5029) in this category, indicating a superior correlation between predictions and ground truth compared to its peers. Interestingly, while CodeT5P leads in F1 and MCC, CodeBERT maintains the highest AUC (0.9018), suggesting a high potential for class discrimination despite its lower precision. Additionally, CodeT5P balanced the trade-off between predicting false positives and false negatives, achieving the lowest FPR (0.0385) among studied PLMs. Although UniXCoder achieves higher Recall (0.8879), indicating sensitivity in identifying vulnerable lines, its significantly lower Precision (0.1787) and higher FPR (0.1120) severely impacts its overall performance, yielding a much lower F1-score of 0.2975. This is also reflected in its lower MCC of 0.3688 compared to CodeT5P. CodeBert exhibits the modest performance among PLMs, with an F1-score of 0.3715, despite having relatively high Recall (0.8822). Thus, among traditional PLMs, CodeT5P emerges as the most viable choice for line-level vulnerability detection, effectively balancing FPR and FNR to mitigate the challenges of the imbalanced dataset.
@@
Observation 6: While using prompting alone, few-shot prompting can improve the effectiveness of LLMs in detecting line-level multilingual vulnerability. Under zero-shot prompting, all evaluated LLMs show limited practical effectiveness despite achieving notably high Accuracy scores. Code Llama attains the highest Accuracy (0.9589); however, it exhibits an extremely low F1-score (0.2254), reflecting significant weaknesses in Precision (0.2272), Recall (0.2237), and FNR (0.7763). The poor discriminative ability is further highlighted by its low MCC (0.2043) and AUC (0.6014). Similarly, models like GPT-4o and GPT-3.5-Turbo display relatively high Accuracies (0.9537 and 0.9557, respectively) but similarly low F1-scores (0.3313 and 0.2797) and high FNR (0.5714 and 0.6785). Their zero-shot MCC values (0.3175 and 0.2596, respectively) and AUC scores (0.6984 and 0.6473) underscore the difficulty these models face in isolating vulnerable lines without specific guidance. The consistently low Recall/Precision and high FNR across zero-shot prompting indicate a clear inability of LLMs to effectively detect vulnerable lines without further guidance or contextual clues, underscoring the insufficiency of using LLMs in isolation under zero-shot conditions for multilingual line-level vulnerability detection.
@@
Few-shot prompting improves LLMs’ performance through in-context examples. GPT-4o demonstrates the best performance, surpassing all other PLMs with a significantly higher F1-score of 0.6050, supported by strong Precision (0.6461) and moderate Recall (0.5689). This improvement is validated by a substantial jump in MCC to 0.5962 and an AUC of 0.7802. Compared to the zero-shot prompting setting, GPT-4o’s FPR significantly decreases from 0.0318 to 0.0086, while its FNR also reduces from 0.5714 to 0.4311. GPT-3.5-Turbo also shows notable improvement, with its F1-score rising to 0.5078. This improvement is primarily driven by enhanced Precision (0.6057), although Recall remains limited (0.4372). Additionally, its MCC increases to 0.5035 and its AUC reaches 0.7147 under the few-shot setting. These results demonstrate that few-shot prompting alone, without instruction tuning, can partially address the issue of imbalanced labels (where vulnerable lines are fewer than non-vulnerable lines) to improve LLMs’ performance in detecting multilingual vulnerabilities at the line-level.
@@
Figure 5: Performance of PLMs and LLMs across seven programming languages in line-level multilingual vulnerability detection (y-axis: F1-score)
@@
Observation 7: Instruction tuning is also beneficial for improving LLMs in detecting multilingual vulnerabilities at the line-level. Finally, the addition of instruction tuning combined with few-shot prompting significantly enhances LLM performance, clearly outperforming all other prompting strategies. GPT-4o, in particular, achieves the highest overall performance with an impressive F1-score of 0.6641. This is further supported by the highest MCC (0.6563) and a strong AUC of 0.8116 among all tested LLM configurations. This high F1-score results from a robust Precision of 0.7012 and substantial Recall of 0.6307, effectively handling the challenge posed by dataset imbalance. Additionally, GPT-4o reaches the lowest FPR of 0.0074 and the lowest FNR of 0.3693 among the studied LLMs (except for Llama 3 with few-shot prompting). This indicates an excellent balance between correctly identifying vulnerable and clean lines. GPT-3.5-Turbo also shows notable improvements under this condition, achieving an F1-score of 0.5215, along with an MCC of 0.5098 and an AUC of 0.7813, demonstrating the broader positive impact of instruction tuning combined with few-shot prompting. While using instruction tuning with few-shot prompting, GPT-4o and GPT-3.5-Turbo significantly outperform the dummy classifier in F1-score and Precision. This demonstrates that their predictions are substantially more accurate than random guesses for line-level detection.
@@
When comparing GPT-4o with the best-performing PLM, CodeT5P (F1-score: 0.4841), GPT-4o’s superiority becomes even more evident. GPT-4o notably surpasses CodeT5P not only in F1-score (0.6641 vs. 0.4841) but also significantly in Precision (0.7012 vs. 0.3536). The superiority of GPT-4o is also reflected in the MCC comparison (0.6563 for GPT-4o vs. 0.5029 for CodeT5P). However, it is worth noting that CodeT5P maintains a higher overall AUC (0.8644) than GPT-4o (0.8116), indicating that while GPT-4o is more effective at the current threshold, the PLM may possess broader discriminative potential across different classification thresholds. Although CodeT5P achieves slightly higher Recall (0.7673 vs. GPT-4o’s 0.6307), its lower Precision substantially limits its practical utility, as it generates many more false-positive predictions. Supporting evidence comes from the FPR and FNR metrics. GPT-4o achieves a much lower FPR (0.0074 vs. CodeT5P’s 0.0385), though it has a higher FNR (0.3693 vs. CodeT5P’s 0.2327).
@@
3.3 RQ3: What are the strengths and weaknesses of the PLMs and LLMs in multilingual vulnerability detection?
@@
  • The unique correct and incorrect detection. To investigate the complementarity of PLMs and LLMs, we performed a set-membership analysis using Venn diagrams. We categorized the results based on the instance-level alignment between model predictions and ground-truth labels. Specifically, we constructed sets of Correct Detections (True Positives and True Negatives) and Incorrect Detections (False Positives and False Negatives) for each model. This approach allows us to quantify the orthogonality of the models—identifying whether LLMs succeed on the specific code instances where PLMs fail, and vice-versa. The Venn diagrams thus represent the distribution of unique and shared performance over the total population of test instances, rather than semantic code features.
@@
Figure 6: The unique correct/incorrect detection of function-level multilingual vulnerability in PLMs and LLMs
Figure 7: The unique correct/incorrect detection of line-level multilingual vulnerability in PLMs and LLMs
@@
Table 5: The percentage of correct detection for function-level multilingual vulnerability detection using two representative techniques across the Top 25 most dangerous CWE-IDs in 2023